There are 850 Douglas fir and ponderosa pine trees in a
section of forest bought by Sawz Logging Co. The company paid an average of
$300 for each Douglas fir and $225 for each ponderosa pine. If the company paid
$217,500 for the trees, how many of each kind did the company buy?

Answer by checkley71(8403)

D+P=850 OR D=850-P
300D+225P=217500 NOW REPLACE D WITH (850-P) & SOLVE FOR P
300(850-P)+225P=217500
255000-300P+225P=217500
-75P=217500-255000
-75P=37500/-75
P=500 NUMBER OF THE PONDEROSA PINES.
850-500=350 NUMBER OF DOUGLAS FIRS.
PROOF
300*350+225*500=217500
105000+112500=217500
217500=217500